Instructions

Please fill the user’s information on the request form, and sign it at the bottom. The user’s supervisor must fill the supervisor’s information and sign the form. In case the user holds a permanent position within the Institution specified on the form, he/she may fill and sign as his/her own supervisor.

Please return the completed form by post to

Alfredo Ferrari EN/STI, Mailbox Z04500 CERN CH-1211 Geneva 23 Switzerland fax: +41-227669474

Sending of a copy by fax may speed up the procedure, however the original form must also be sent by standard mail.